# On Resolution of $\ell_{1}$-norm minimization via a Two-metric Adaptive Projection Method

### v1

@@ 6 Discussion
Unfortunately, it will lead to numerical instability if we directly apply the two-metric projection method to Eq. 32. To illustrate this, we consider the eigendecomposition of Hessian.
$\nabla^{2}f(x_{k})=Q^{T}\Lambda Q$
Let $Q=[q_{1},q_{2},\ldots,q_{n}]$, and $Q_{1}=[q_{i_{1}}]_{i_{1}\in I_{k}^{1}}$, $Q_{2}=[q_{i_{2}}]_{i_{2}\in I_{k}^{2}}$. Then, in two-metric projection, we need to solve a Newton equation as follows,
@@
Another salvage is using (limited) BFGS to approximate Hessian. Schmidt proposed ProjectionL1 for problem (32) in [20], and this is a variant of the two-metric projection method that we discuss in this section. They used two Quasi-Newton methods to approximate Hessian: BFGS and L-BFGS. L-BFGS is more suitable to large-scale problems.
@@ 7 Conclusion
In this paper, we propose a two-metric adaptive projection method for solving the $\ell_{1}$-norm minimization problem. Compared with the two-metric projection method proposed by Bertsekas and ProjectionL1 proposed by Schmidt to solve the equivalent bound-constrained problem, our algorithm does not require an exact solver of Newton system or the strong convexity of $f$. Moreover, it can be shown to converge globally to an optimal solution and have local superlinear convergence rate. The key to our analysis is the Error Bound condition, which is a weaker condition than the strong convexity of $f$ and sufficient second-order condition for problem Eq. 1. We also provide a complexity guarantee for our algorithm in the nonconvex setting.
@@
There are several future directions worth studying on the two-metric adaptive projection method. First, we can extend our algorithm to solve other separable regularized optimization problems, such as the $\ell_{0}$-regularized problem. Second, we can use LBFGS to approximate the Hessian matrix in our algorithm to improve the computational efficiency.
@@ Acknowledgments
We would like to acknowledge the valuable suggestions provided by Professor Stephen J. Wright when writing this manuscript.

### v2

@@ 6 Discussion
Unfortunately, it will lead to numerical instability if we directly apply the two-metric projection method to Eq. 33. To illustrate this, we consider the following eigendecomposition of Hessian,
$\nabla^{2}f(x_{k})=Q^{T}\Lambda Q.$
Let $Q\triangleq[q_{1},q_{2},\ldots,q_{n}]$, and $Q_{1}\triangleq[q_{i_{1}}]_{i_{1}\in I_{k}^{1}}$, $Q_{2}\triangleq[q_{i_{2}}]_{i_{2}\in I_{k}^{2}}$. Then, in two-metric projection, we need to solve a Newton equation as follows,
@@
Another salvage is using (limited) BFGS to approximate Hessian. Schmidt proposed ProjectionL1 for problem (33) in [25], and this is a variant of the two-metric projection method that we discuss in this section. They used two Quasi-Newton methods to approximate Hessian: BFGS and L-BFGS. L-BFGS is more suitable for large-scale problems.
@@ 7 Conclusion
In this paper, we propose a two-metric adaptive projection method for solving the $\ell_{1}$-norm minimization problem. Compared with the two-metric projection method proposed by Bertsekas and ProjectionL1 proposed by Schmidt to solve the equivalent bound-constrained problem, our algorithm does not require an exact solver of Newton system or the strong convexity of $f$. Moreover, it can be shown to converge globally to an optimal solution and have manifold identification property as well as local superlinear convergence rate. One of the keys to our analysis is the Error Bound condition, which is a weaker condition than the strong convexity of $f$ and sufficient second-order condition for problem Eq. 1.
@@
There are several future directions worth studying on the two-metric adaptive projection method. First, we can extend our algorithm to solve other separable regularized optimization problems, such as the $\ell_{0}$-regularized problem. Second, we can consider quasi-Newton methods (such as LBFGS) to potentially further improve the computational efficiency.
@@ Acknowledgments
We would like to acknowledge the valuable suggestions provided by Professor Stephen J. Wright, Professor Kim-Chuan Toh and Professor Chingpei Lee when writing this manuscript.